OpenStudy (anonymous):

how many solutions are there for 8x^2 - 13x + 3 = 0

hero (hero):

b^2 - 4ac = 13^2 - 4(8)(3) = 169 - 96 = 73 2 solutions

OpenStudy (anonymous):

ok thanks i have another question

OpenStudy (anonymous):

The perimeter of a rectangle is 54 cm. The area of the same rectangle is 176 cm². What are the dimensions of the rectangle?

hero (hero):

P = 2(l+w) A = lw 54 = 2(l + w) 176 = lw 54 = 2l + 2w 176 = lw Solve the system

hero (hero):

l = (54 - 2w)/2 l = 176/w l = l w(54 - 2w) = 2(176) 54w - 2w^2 = 352 2w^2 - 54w + 352 = 0 w^2 - 27w + 176 = 0 w^2 - 11w - 16w + 176 = 0 w(w-11) - 16(w -11) = 0 (w-11)(w-16) = 0 w = 11 l = 16

hero (hero):

Re-post as a separate question.